# When Knockoffs fail: diagnosing and fixing non-exchangeability of Knockoffs

### v1

@@ 3.1 Gaussian Knockoffs
Candès et al., 2018 have built Knockoffs that are provably exchangeable for Gaussian data. Provided an observed $X$ with $X\sim\mathcal{N}(0,\Sigma)$, we sample from
$\tilde{X}\mid X\stackrel{{\scriptstyle\mathrm{d}}}{{=}}\mathcal{N}(\mu,\mathbf{V}),$
where
$\begin{gathered}\mu=X-X\boldsymbol{\Sigma}^{-1}\operatorname{diag}(s)\\ \mathbf{V}=2\operatorname{diag}(s)-\operatorname{diag}(s)\boldsymbol{\Sigma}^{-1}\operatorname{diag}(s)\end{gathered}$
@@
With $\operatorname{diag}(s)$ any diagonal matrix. Then, if $\operatorname{diag}(s)$ is chosen such that $\mathbf{G}$ is positive semidefinite, we obtain valid Knockoffs since:
@@
$[X,\tilde{X}]\sim\mathcal{N}(0,\mathbf{G}),\quad\text{ where }\mathbf{G}=\left(\begin{array}[]{cc}\boldsymbol{\Sigma}&\boldsymbol{\Sigma}-\operatorname{diag}(s)\\ \boldsymbol{\Sigma}-\operatorname{diag}(s)&\boldsymbol{\Sigma}\end{array}\right)$
A major challenge is that the covariance matrix $\boldsymbol{\Sigma}$ is unknown and has to be estimated via shrinkage procedures such as Ledoit-Wolf shrinkage (Ledoit and Wolf, 2003) or Graphical Lasso (Friedman, Hastie and Tibshirani, 2008). Given a correctly estimated $\boldsymbol{\Sigma}$, examples of valid $s$ include the equicorrelated construction $s_{j}^{\mathrm{EQ}}=2\lambda_{\min}(\Sigma)\wedge 1$ for all $j$.
@@
In practice, testing exchangeability for non-Gaussian data is challenging. In the case of Gaussian data, since the actual covariance is not known but has to be estimated, exchangeability depends directly on the accuracy of the covariance estimate. Even if the data is Gaussian, its covariance matrix may be hard to estimate properly in high-dimensional regimes (Stein, Efron and Morris, 1972; Fukunaga, 2013). Indeed, such covariance estimation has to draw a difficult compromise between data fit and positive definiteness, along sparsity of the inverse covariance estimation from the Graphical Lasso. L2 shrinkage covariance estimation (Ledoit and Wolf, 2003) on the other hand, is known to lead to excessive bias (Belilovsky et al., 2017). In addition to the sheer computation cost, hyperparameter setting for the Graphical Lasso is challenging, and leads to difficult and costly parameter selection.
3.2 Existing alternatives to Gaussian Knockoffs
Other available methods to build Knockoffs include Monte-Carlo based methods (Sesia, Sabatti and Candès, 2019; Bates et al., 2021). These procedures focus on discrete distributions as in the case of GWAS studies. In the continuous setting, the method proposed in Bates et al., 2021 is equivalent to Gaussian Knockoffs.
@@ 3.2 Existing alternatives to Gaussian Knockoffs
Recent machine learning techniques such as deep learning have also been used to apprehend complex data dependence structures in the context of Knockoffs. Romano, Sesia and Candès, 2020 propose using Deep Neural networks to build Knockoffs using a target covariance matrix estimated from the data as in the Gaussian algorithm. Additionally, the proposed algorithm aims at minimizing the correlation between original variables and Knockoffs to maximize power. Using similar ideas, Zhu et al., 2021; Liu and Zheng, 2018 propose using a Variational Auto-Encoders (VAE) to build Knockoffs as VAEs allow learning and sampling from complex data distributions. However, in the context of high-dimensional variable selection, the number of samples available may be insufficient to properly train a Deep Learning model. A limitation of these approaches is that they offer no theoretical guarantees on Knockoffs exchangeability.
3.3 An efficient Non-parametric Knockoff generation algorithm
We propose an alternative method to build Knockoffs by learning to predict $X_{j}$ from $X_{-j}$. This idea is related to the Sequential Conditional Independent Pairs algorithm of Candès et al. (2018). A related idea is exploited to quantify variable importance in Random Forests in Chamma, Engemann and Thirion, 2023. The proposed approach is described in Algorithm 1.
In this approach, the dependency structure of the data is learned using a machine learning model $f$ for each variable that predicts $X_{j}$ from $X_{-j}$. Typically, using a Lasso model is a good default choice. In practice, we set $\lambda=\lambda_{max}/100$ with $\lambda_{max}=\frac{1}{n}\left\|\mathbf{X_{-j}}^{T}X_{j}\right\|_{\infty}$. The $j$-th fitted predictor is denoted $f_{j}$. Then, the residual $\widehat{\epsilon}_{j}=X_{j}-f_{j}(\mathbf{X_{-j}})$ is computed. Once this is done for all variables, Knockoffs are built by shuffling the residuals: $\tilde{X}_{j}\text{ is chosen as }f_{j}(\mathbf{X_{-j}})+\widehat{\epsilon}_{\tau{(j)}}$ with $\tau$ a permutation of $\llbracket p\rrbracket$. The important point is that this approach circumvents the difficult problem of covariance estimation in high-dimensional settings. Note that using a Lasso model is suitable in regimes where $n\simeq p$ – if $n\gg p$ selecting a more expressive model such as random forests (Breiman, 2001) or gradient boosting trees (Chen et al., 2015) is necessary to retain control.
@@ 3.3 An efficient Non-parametric Knockoff generation algorithm
2      Fit a prediction model $f_{j}$ on $(\mathbf{X_{-j}},X_{j})$ // Typically a Lasso model
3       Compute the residual $\widehat{\epsilon}_{j}=X_{j}-f_{j}(\mathbf{X_{-j}})$
@@
7      Sample $\text{Sample }\tilde{X}_{j}=f_{j}(\mathbf{X_{-j}})+\widehat{\epsilon}_{\tau{(j)}}$ // $\tau$ is a permutation of $\llbracket p\rrbracket$

### v2

@@ 3.1 Gaussian knockoffs
Candès et al.,, 2018 have introduced a knockoff construction framework for Gaussian data. Given $X\sim\mathcal{N}(0,\boldsymbol{\Sigma})$, we sample $\tilde{X}$ from $\mathcal{N}(\mu,\mathbf{V})$, where $\mu=X-X\boldsymbol{\Sigma}^{-1}\operatorname{diag}(\mathbf{s})$ and $\mathbf{V}=2\operatorname{diag}(\mathbf{s})-\operatorname{diag}(\mathbf{s})\boldsymbol{\Sigma}^{-1}\operatorname{diag}(\mathbf{s})$, with $\operatorname{diag}(\mathbf{s})$ any diagonal matrix. If $\operatorname{diag}(\mathbf{s})$ is chosen such that $\mathbf{G}$ is positive semidefinite, we obtain valid knockoffs since:
@@
$[X,\tilde{X}]\sim\mathcal{N}(0,\mathbf{G})\quad\text{where}\quad\mathbf{G}=\left(\begin{array}[]{cc}\boldsymbol{\Sigma}&\boldsymbol{\Sigma}-\operatorname{diag}(\mathbf{s})\\ \boldsymbol{\Sigma}-\operatorname{diag}(\mathbf{s})&\boldsymbol{\Sigma}\end{array}\right).$
In particular, this is the case with the equi-correlated construction where $s_{j}^{\mathrm{EQ}}=2\lambda_{\min}(\boldsymbol{\Sigma})\wedge\frac{1}{p}tr(\boldsymbol{\Sigma})$ for all $j$.
@@
In practice, even when $X$ can be assumed to be Gaussian, its covariance matrix $\boldsymbol{\Sigma}$ is unknown and must be estimated from the data. This is generally done via shrinkage procedures such as Ledoit-Wolf shrinkage (Ledoit and Wolf,, 2003) or Graphical Lasso (Friedman et al.,, 2008). The validity of the resulting knockoffs directly depends on the accuracy of the covariance estimate. Covariance estimation in high-dimensional regimes is a recognized as a difficult problem (Stein et al.,, 1972; Fukunaga,, 2013). Indeed, this estimation requires to find an acceptable compromise between data fit and positive definiteness, along sparsity of the inverse covariance estimation from the Graphical Lasso. L2 shrinkage covariance estimation (Ledoit and Wolf,, 2003) on the other hand, is known to lead to excessive bias (Belilovsky et al.,, 2017). In addition to the intrinsic computation cost, hyperparameter setting for the Graphical Lasso is challenging, and leads to difficult and costly parameter selection.
3.2 Existing nonparametric knockoff generation methods
Recent machine learning techniques, such as deep learning, have also been used to capture complex data dependence structures in the context of knockoffs. Romano et al., (2020) propose using deep neural networks to build knockoffs based on a target covariance matrix estimated from the data, as in the Gaussian algorithm. This approach represents a second-order approximation. The loss function also includes a term derived from the Maximum Mean Discrepancy to account for higher-order effects.
Furthermore, to ensure a powerful procedure, the knockoff covariates should be as different as possible from the original ones. To achieve this, an additional term is incorporated to penalize high correlation between them.
@@ 3.2 Existing nonparametric knockoff generation methods
Using similar ideas, Zhu et al., (2021) and Liu and Zheng, (2018) propose using a Variational Auto-Encoder (VAE) to build knockoffs as VAEs allow learning and sampling from complex data distributions. However, in the context of high-dimensional variable selection, the number of available samples may be insufficient to properly train a Deep Learning model. A limitation of these approaches is that they offer no theoretical guarantees on knockoffs exchangeability.
3.3 Knockoff generation methods derived from the SCIP algorithm
3.3.1 The Sequential Conditional Independent Pairs algorithm
Candès et al.,, 2018 have introduced a theoretical nonparametric knockoff generation approach known as the Sequential Conditional Independent Pairs (SCIP) algorithm. It consists in progressively learning the distribution $X_{j}$ given $\mathbf{x}_{-j}$ and the previous knockoffs as described in Algorithm 1.
$j=1$
2 while $j\leq p$ do
3       Sample $\tilde{X}_{j}$ from $\mathcal{L}\left(X_{j}\mid X_{-j},\tilde{X}_{1:j-1}\right)$
4       $j=j+1$
5 end while
Algorithm 1 Sequential Conditional Independent Pairs (SCIP; Candès et al.,, 2018)
This algorithm builds valid knockoffs assuming all the conditional distributions $\mathcal{L}\left(X_{j}\mid X_{-j},\tilde{X}_{1:j-1}\right)$ to be known (Candès et al.,, 2018). However, as also noted by Candès et al., (2018), this knowledge is generally not available in practice. Another caveat of this approach is its sequential nature: Algorithm 1 requires computing a new conditional distribution at each step without the possibility of parallelization, making computations potentially intractable. These two practical limitations are discussed and addressed in the remainder of this section.
3.3.2 Learning the conditional distributions
To instantiate Algorithm 1, we propose using a machine learning model $f\in\mathcal{F}$ to learn the conditional distributions. A related idea is used for Variable Importance in the Conditional Permutation Importance of Chamma et al.,, 2023 and proved valid in Reyero Lobo et al.,, 2025. We observe from Algorithm 1 that the order in which the knockoff variables are constructed is arbitrary. Therefore, for simplicity, we fix the order to match the input order. For each $j$, $f_{j}$ predicts $X_{j}$ from $X_{-j},\tilde{X}_{1:j-1}$. Typically, using a Lasso model is a good default choice because, in the framework of Candès et al.,, 2018, the relationship between the input covariates is assumed to be simple, and this choice helps to keep the algorithm computationally tractable. In practice, we set $\lambda=\lambda_{max}/100$ with $\lambda_{max}=\frac{1}{n}\left\|(\mathbf{X}_{-j},\mathbf{\tilde{X}}_{1:j-1})^{T}\mathbf{X}_{j}\right\|_{\infty}$. Then, the residual $\widehat{\boldsymbol{\epsilon}}_{j}=\mathbf{X}_{j}-f_{j}(\mathbf{X}_{-j},\mathbf{\tilde{X}}_{1:j-1})$ is computed. The $j$-th knockoff variable is built by drawing a residual at random: $\tilde{\mathbf{X}}_{j}=f_{j}(\mathbf{X_{-j},\tilde{X}_{1:j-1}})+\sigma({\widehat{\boldsymbol{\epsilon}}_{j}})$, where $\sigma$ is a permutation of $\llbracket n\rrbracket$. The complete approach is described in Algorithm 2.
Input : $f\in\mathcal{F}$
1 for $j\in[1,p]$ do
2      Fit a prediction model $f_{j}$ on $((\mathbf{X}_{-j},\mathbf{\tilde{X}}_{1:j-1}),\mathbf{X}_{j})$ // Typically a Lasso model
3       Compute the residual $\widehat{\boldsymbol{\epsilon}}_{j}=\mathbf{X}_{j}-f_{j}((\mathbf{X}_{-j},\mathbf{\tilde{X}}_{1:j-1}))$
4       Sample $\text{Sample }\mathbf{\tilde{X}}_{j}=f_{j}((\mathbf{X}_{-j},\mathbf{\tilde{X}}_{1:j-1}))+\sigma(\widehat{\boldsymbol{\epsilon}}_{j})$
5 end for
Return $\mathbf{\tilde{X}_{1:p}}$
Algorithm 2 Sequential generation of nonparametric knockoffs by learning to predict $\mathbf{X}_{j}$ from $(\mathbf{X_{-j},\tilde{X}_{1:j-1}})$ using a model $f_{j}$.
We turn to showing that under mild assumptions, this indeed amounts to sampling from $\mathcal{L}\left(X_{j}\mid X_{-j},\tilde{X}_{1:j-1}\right)$. In practice, we need to take into account the fact that $f^{*}_{j}\left(X_{-j},\tilde{X}_{1:j-1}\right):=\mathbb{E}\left[X_{j}\mid X_{-j},\tilde{X}_{1:j-1}\right]$ is estimated from a finite number of samples. To ensure asymptotic guarantees on the generated distribution, we require each regressor to be sufficiently accurate:
Definition 3 (Estimator consistency)
Given a training set of size $n$ used to regress an output $Y$ given an input $X$, a regressor $f_{n}$ is said to be consistent if
$\mathbb{E}\left[\left(f_{n}(X)-\mathbb{E}\left[Y|X\right]\right)^{2}\right]\xrightarrow[\enskip n\to\infty\enskip]{}0.$
In the following result, we establish the asymptotic validity the above sampling scheme: under the assumption of consistent regressors, the Wasserstein-2 distance between the samples produced by Algorithm 2 and the desired conditional distribution converges to 0 as the sample size grows to infinity.
Proposition 1 (Validity of auto-regressive sampling)
For each $j\in\llbracket p\rrbracket$, assuming that $X$ is Gaussian and assuming the consistency of each regressor $f_{j}\in\mathcal{F}$, with $\mathcal{F}$ a $P$-Donsker class, then $\tilde{X}_{j}|X_{-j},\tilde{X}_{1:j-1}\sim\widehat{P}_{j}$ such that $\widehat{P}_{j}\xrightarrow[n\to\infty]{W_{2}}\mathcal{L}\left(X_{j}\mid X_{-j},\tilde{X}_{1:j-1}\right)$.
3.3.3 Parallel knockoff generation
In practice, Algorithm 2 is computationally costly to run. Fitting the models $f_{j}$ cannot be parallelized as each fit depends on previously built knockoffs. To make computations tractable, we propose an approximate version of this Algorithm, in which samples are drawn from $\mathcal{L}\left(X_{j}\mid X_{-j}\right)$ instead of $\mathcal{L}\left(X_{j}\mid X_{-j},\tilde{X}_{1:j-1}\right)$. In practice, this allows us to fit all models $f_{j}$ in parallel. Once this is done for all variables, knockoffs are built by shuffling the residuals: $\mathbf{\tilde{X}}_{j}\text{ is chosen as }f_{j}(\mathbf{X}_{-j})+\sigma(\widehat{\boldsymbol{\epsilon}}_{j})$ with $\sigma$ a permutation of $\llbracket n\rrbracket$.
In words, we remove conditioning on previously built knockoff variables to make the algorithm easily parallelizable, as described in Algorithm 3.
@@ 3.3.3 Parallel knockoff generation
2      Fit a prediction model $f_{j}$ on $(\mathbf{X}_{-j},\mathbf{X}_{j})$ // Typically a Lasso model
3       Compute the residual $\widehat{\boldsymbol{\epsilon}}_{j}=\mathbf{X}_{j}-f_{j}(\mathbf{X}_{-j})$
@@
7      Sample $\text{Sample }\mathbf{\tilde{X}}_{j}=f_{j}(\mathbf{X_{-j}})+\sigma({\widehat{\boldsymbol{\epsilon}}_{j}})$ // $\sigma$ is a permutation of $\llbracket n\rrbracket$